# On the Generalization of Preference Learning with DPO

### v3

@@ Acknowledgement
as $l_{b}\leq 1$.
Appendix B Multi-Token Derivation
@@
We start from the Equation (14),
@@
$-\beta\sum_{i=1}^{N}\sigma\big{(}r(y_{l,i})-r(y_{w,i})\big{)}\sum_{j=1}^{L}\frac{\partial\log\mathcal{S}(Wg(i,j,w))}{\partial W}-\frac{\partial\log\mathcal{S}(Wg(i,j,l))^{\top}\mathbf{y}_{w/l,i}^{(j)}}{\partial W}$ (57)
@@
Appendix C Future Extension Beyond DPO
Our work focuses on reward generalization behavior for preference learning specifically for DPO, but the framework presented can be extended to a more general class of objectives, in particular, the family of objectives presented in GPO [58] and also SimPO [59] with fixed length responses. This is because the objective function is of the form,
$\mathcal{L}(\pi_{\theta};\pi_{\text{ref}};\mathcal{D})=-\mathbb{E}_{(x,y_{w},y_{l})\in\mathcal{D}}\bigg{[}f\bigg{(}\beta\bigg{(}\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\bigg{)}\bigg{)}\bigg{]},$ (60)
and the only modification to the dynamics would be replacing the $\sigma(-r_{i})$ factor in
with $f^{\prime}(r_{i})$. This points towards a promising direction of developing conditions under which the behavior of other preference learning methods can be guaranteed. We leave this as future work.
Appendix D Additional Verification
@@
Here we provide the plot of the cosine similarities of embeddings between different personas before and after subtracting the mean embedding in Figure 5(a) and 5(b). The personas are ordered according to lexicographical order.
@@
We present the training and test losses and accuracies across different numbers of clusters as seen in Figures 6(a), 6(b), 7(a), and 7(b). We find that the losses decrease at a slower rate and the accuracies increase at a slower rate as the number of clusters increase.
@@
Verification on Llama-3.1-8B
We provide verification of the generalization results with the same training setup as with LlaMa-2-7B and provide the results in Figures 8(a), 8(b), 9(a), 9(b), 10(a), 10(b).
Figure 8: Llama-3.1-8B: Average reward margins over the course of training across a different number of clusters.
Figure 9: Llama-3.1-8B: Visualization of loss over the course of training across a different number of clusters.
Figure 10: Llama-3.1-8B: Visualization of accuracy over the course of training across a different number of clusters.
Appendix E Training and Experimental Details
@@
For all training runs, we use the AdamW optimizer with a learning rate of 1e-5 with no warm-up steps and a constant learning rate. We train on 4 GPUs with a batch size of 32 per device.
@@ 2 Clusters: desire-for-no-human-oversight-sometimes, agreeableness
We train with 4 A100 80GB GPUs using the TRL library [89] and Huggingface library [90], generate embeddings with the Huggingface library and 1 A100 80GB GPU, and perform all other parts of the experiments on a AMD EPYC 7513 32-Core Processor CPU. The total time to reproduce all experiments is estimated to be 6 hours.

### v5

@@ Conclusion.
as $l_{b}\leq\frac{1}{2}$.
Appendix C Approximate-orthogonal clusters
In this section, we prove extensions of Theorems 4.2 and 4.3 under approximately orthogonal clusters.
We start with our definition of $\delta$-approximately orthogonal clusters.
Definition C.1 ($\delta$-approximately orthogonal clusters).
We consider clusters distributed as $\mathcal{N}(l_{b}e_{1}\pm c_{i},v^{2}I_{d})$ where each $c_{i}$ has a corresponding standard basis vector $e_{c_{i}}$. Then, we will define the clusters as being $\delta$-approximately orthogonal with $0\leq\delta<\frac{1}{2}$ if for every $i$, $|\langle c_{i},e_{c_{i}}\rangle|\geq 1-\delta^{2}/8$.
Lemma C.2 (Pairwise approximate orthogonality).
Given $\delta$-approximately orthogonal clusters and $i_{1},i_{2}$ from $[K]$ with $i_{1}\neq i_{2}$, then we have that
$|\langle c_{i_{1}},c_{i_{2}}\rangle|\leq\delta$ (48)
Proof.
Without loss of generality, assume that the dot products of $c_{i_{1}},c_{i_{2}}$ and their corresponding standard basis vectors are positive. Let $\xi_{1}$ be the angle between $c_{i_{1}}$ and $e_{c_{i_{1}}}$, and let $\xi_{2}$ be the angle between $c_{i_{2}}$ and $e_{c_{i_{2}}}$. Then, by $\delta$-approximate orthogonality, we have that $\xi_{1},\xi_{2}\leq\arccos(1-\delta^{2}/8)$. Then, in the worst case, we have that
$|\langle c_{i_{1}},c_{i_{2}}\rangle|=\sin(2\arccos(1-\delta^{2}/8))=2(1-\delta^{2}/8)\sqrt{1-(1-\delta^{2}/8)^{2}}$ (49)
Then,
$|\langle c_{i_{1}},c_{i_{2}}\rangle|\leq 2\sqrt{\delta^{2}/4-\delta^{4}/64}\leq\delta$ (50)
Lemma C.3.
Given $\delta$-approximately orthogonal clusters with $\delta\leq\min\left(\frac{\epsilon v}{8},\frac{1}{4}\right)$, $l_{b}\leq\frac{1}{2}$, then with probability at least $1-(16Z+8)KQ^{2}e^{-\epsilon^{2}/16}-(8Z+4)KQ^{2}\exp\left(-\frac{c\epsilon}{v}\min\left(1,\frac{\epsilon}{dv}\right)\right)$, for any $i\in[K]$ and for any $j,k\in[Q]$
for any $i\in[K]$ and for any $j,k\in[Q]$
for any $i_{1}\neq i_{2}$ that share a token and for any $j,k\in[Q]$
$\left|C(x_{j}^{(i_{1},\pm)},x_{k}^{(i_{2},\pm)})\right|\leq l_{b}^{2}+\frac{5}{2}\epsilon v$ (54)
$\left|C(x_{j}^{(i_{1},\pm)},x_{k}^{(i_{2},\mp)})\right|\leq l_{b}^{2}+\frac{5}{2}\epsilon v$ (55)
Proof.
For each dot product between a pair of inputs from the same set of clusters, we introduce an additional deviation of at most $l_{b}\delta+\delta\alpha$ from the dot product between the deviation of $c_{i}$ from $e_{c_{i}}$ and $e_{1}$, the dot product between the deviation of $c_{i}$ from $e_{c_{i}}$ and the Gaussian component. The contribution from the dot product between $c_{i}$’s does not change in this case. Since $\delta\leq\frac{\epsilon v}{8}$ and $\delta\leq\frac{1}{4}$, we can use the same bounds on the $\alpha$’s so that with probability at least $1-8KQ^{2}e^{-\epsilon^{2}/16}$, we have
for any $i\in[K]$ and for any $j,k\in[Q]$
Now, we consider (54). We know that in the case of Lemma B.1, we have
In this case, we have additional terms coming due to approximate-orthogonality. We bound the difference between the right hand side above and the current $x_{j}^{(i_{1},\pm)}\cdot x_{k}^{(i_{2},\pm)}$. In particular, we have most $2\delta$ from the dot products between $c_{i_{1}},c_{i_{2}}$ and at most $\delta l_{b}$ from the dot products between $c_{i,1}$ or $c_{i_{2}}$ and $e_{1}$. We have an additional shift of at most $\frac{\delta}{2}(\alpha_{j,e_{d_{1}}}+\alpha_{k,e_{d_{2}}})$ for some $d_{1},d_{2}$ from the a change of at most $\delta/2$ along a new basis vector. Then, we introduce a change of at most
$2\delta+\delta l_{b}+\frac{\delta}{2}(\alpha_{j,e_{d_{1}}}+\alpha_{k,e_{d_{2}}})$ (59)
which in the worst case for all $i_{1},i_{2},j,k$ is at most with probability at most $1-8ZKQ^{2}e^{-\epsilon^{2}/16}$
$\frac{\epsilon v}{2}$ (60)
Then, we have that
$\left|C(x_{j}^{(i,\pm)},x_{k}^{(i,\pm)})\right|\leq l_{b}^{2}+\frac{5}{2}\epsilon v$
(13) follows similarly. Then, the full result holds with probability at least $1-(16Z+8)KQ^{2}e^{-\epsilon^{2}/16}-(8Z+4)KQ^{2}\exp\left(-\frac{c\epsilon}{v}\min\left(1,\frac{\epsilon}{dv}\right)\right)$ for some constant $c>0$.
Lemma C.4.
With $\delta\leq\min\left(\frac{\epsilon v}{8},\frac{1}{4}\right),l_{b}\leq 1$, with probability at least $1-(16Z+8)KQ^{2}e^{-\epsilon^{2}/16}-(8Z+4)KQ^{2}\exp\left(-\frac{c\epsilon}{v}\min\left(1,\frac{\epsilon}{dv}\right)\right)$, we have that for each sample,
$\left|\tau\dot{r_{j}}^{i,\pm}-\frac{2(1+l_{b}^{2})\beta^{2}}{N}\sum_{m=1}^{Q}\sigma(-r_{m}^{i,\pm})-\frac{2(1-l_{b}^{2})\beta^{2}}{N}\sum_{m=1}^{Q}\sigma(-r_{m}^{i,\mp})-\frac{2dv^{2}\beta^{2}}{N}\sigma(-r_{j}^{i,\pm})\right|\\ \leq\frac{2\beta^{2}Q}{N}\left(\left(\frac{5}{2}Z+5\right)\epsilon v+l_{b}^{2}Z\right)\max_{j\in N}\sigma(-r_{j})$ (61)
Proof.
The proof follow exactly the same argument as Lemma B.2.
Theorem C.5.
Given $Z\leq\frac{1}{4l_{b}^{2}}$, $d\leq 5Q$, $v\leq\frac{1}{4\sqrt{Q}}$, and $\epsilon\leq\frac{1}{16v(Z+2)}$, $\frac{1}{2}\geq l_{b}\geq\frac{1}{4}$, $\delta\leq\min\left(\frac{\epsilon v}{8},\frac{1}{4}\right)$, with probability at least $1-(16Z+8)KQ^{2}e^{-\epsilon^{2}/16}-(8Z+4)KQ^{2}\exp\left(-\frac{c\epsilon}{v}\min\left(1,\frac{\epsilon}{dv}\right)\right)$, the trajectory $r_{i}(t)$ for all $i\in[N]$ is upper bounded by $r^{U}(t)$ and lower bounded by $r^{L}(t)$ which are given by
$r^{L}(t)=\frac{Q\beta^{2}}{8N\tau}t$
$r^{U}(t)=\frac{10Q\beta^{2}}{N\tau}t$
for $t\leq\tau_{1}$ and $\tau_{1}$ is given by
$\tau_{1}=\frac{N\tau\log 3}{10Q\beta^{2}}$ (62)
and at $\tau_{1}$, for any training sample $\frac{\log 3}{80}\leq r(t)\leq\log 3$.
Proof.
The proof follows the same argument as Theorem B.3 except now since the right hand side of Lemma C.4 is increased by a factor of $5/4$, for $r^{L}$, the factor of $1/4$ which came from a lower bound $\frac{1}{2}-2(\frac{1}{8})$, is now lower bounded by a factor of $1/8$ as $\frac{1}{2}-\frac{5}{2}(\frac{1}{8})\geq\frac{1}{8}$. For $r^{U}$ this change has no effect.
Theorem C.6.
Given $Z\leq\frac{1}{4l_{b}^{2}}$, $d\leq 5Q$, $v\leq\frac{1}{4\sqrt{Q}}$, and $\epsilon\leq\frac{1}{16v(Z+2)}$, $\delta\leq\min\left(\frac{\epsilon v}{8},\frac{1}{4}\right)$ $\frac{1}{2}\geq l_{b}\geq\frac{1}{4}$, with probability at least $1-(16Z+8)KQ^{2}e^{-\epsilon^{2}/16}-(8Z+4)KQ^{2}\exp\left(-\frac{c\epsilon}{v}\min\left(1,\frac{\epsilon}{dv}\right)\right)$, the generalization error of the implicit reward model at $\tau_{1}$ is bounded as
Proof.
[…]
Then, we will condition on the training set and on the event that Lemma C.2 holds. Then, from Lemma C.2, we know that
$\sum_{m=1}^{d}\alpha_{k,m}^{2}\leq dv^{2}+\epsilon v$ (67)
and we also have that
$|\mu^{(\tilde{i})\top}x^{(\tilde{i})}_{k}-(1+l_{b}^{2})|\leq\frac{5}{2}\epsilon v$
$|\mu^{(\tilde{i})\top}x^{(-\tilde{i})}_{k}-(l_{b}^{2}-1)|\leq\frac{5}{2}\epsilon v$
$|\mu^{(\tilde{i})\top}x^{(j,\pm)}_{k}-l_{b}^{2}|\leq\frac{5}{2}\epsilon v$
Then, $(\tilde{x}-\mu_{\tilde{i}}\top)x_{j}$ conditioned on $x_{j}$ is a centered normal random variable with variance at most $(1+l_{b}^{2}+dv^{2}+\frac{5}{2}\epsilon v)v^{2}$. Then we have that for $\tilde{x}$ with probability at least $1-2KQ^{2}e^{-\epsilon^{2}/2(1+l_{b}^{2}+dv^{2}+\frac{5}{2}\epsilon v)}$ conditioned on the event that Lemma A.1 holds that for any $k\in[Q]$
and for any $i_{2}\in S_{\tilde{i}}$ and for any $k\in[Q]$
$\left|C(\tilde{x},x_{k}^{(i_{2},\pm)})\right|\leq l_{b}^{2}+\frac{7}{2}\epsilon v$ (70)
$\left|C(\tilde{x},x_{k}^{(i_{2},\mp)})\right|\leq l_{b}^{2}+\frac{7}{2}\epsilon v$ (71)
We will condition on the event that the above holds for the remainder of the proof. Then, we have that by the same arguments as in Lemma C.3 that
$\left|\tau\dot{\tilde{r}}-\frac{2(1+l_{b}^{2})\beta^{2}}{N}\sum_{m=1}^{Q}\sigma(-r_{m}^{\tilde{i},\pm})-\frac{2(1-l_{b}^{2})\beta^{2}}{N}\sum_{m=1}^{Q}\sigma(-r_{m}^{\tilde{i},\mp})\right|\\ \leq\frac{2\beta^{2}Q}{N}\left((\frac{7}{2}Z+7)\epsilon v+l_{b}^{2}Z\right)\max_{j\in N}\sigma(-r_{j})$ (72)
and we that that $\tau\dot{\tilde{r}}$ is lower bounded by
$\frac{2(1+l_{b}^{2})\beta^{2}}{N}\sum_{m=1}^{Q}\sigma(-r_{m}^{\tilde{i},\pm})-\frac{2(1-l_{b}^{2})\beta^{2}}{N}\sum_{m=1}^{Q}\sigma(-r_{m}^{\tilde{i},\mp})\\ -\frac{2\beta^{2}Q}{N}\left((\frac{7}{2}Z+7)\epsilon v+l_{b}^{2}Z\right)\max_{j\in N}\sigma(-r_{j})$ (73)
and for $t\leq\tau_{1}$, this is lower bounded by
$\frac{Q\beta^{2}}{N}-\frac{\beta^{2}Q}{N}\left((\frac{7}{2}Z+7)\epsilon v+l_{b}^{2}Z\right)$ (74)
as we know for any training sample $0\leq r_{j}\leq\log 3$. Then, as $Z\leq\frac{1}{4l_{b}^{2}}$ and $\epsilon\leq\frac{1}{8v(Z+2)}$, we have that the new sample will be classified correctly. Then we have that with probability at least $1-(16Z+8)KQ^{2}e^{-\epsilon^{2}/16}-(8Z+4)KQ^{2}\exp\left(-\frac{c\epsilon}{v}\min\left(1,\frac{\epsilon}{dv}\right)\right)$,
as $l_{b}\leq\frac{1}{2}$.
Appendix D Multi-Token Derivation
@@
We start from the equation,
@@
$\beta\sum_{i=1}^{N}\sigma\big(r(y_{l,i})-r(y_{w,i})\big)\sum_{j=1}^{L}\frac{\partial\log\mathcal{S}(Wg(i,j,w))}{\partial W}-\frac{\partial\log\mathcal{S}(Wg(i,j,l))^{\top}\mathbf{y}_{w/l,i}^{(j)}}{\partial W}$ (78)
@@
Appendix E Additional Verification
@@
Here we provide the plot of the cosine similarities of embeddings between different personas before and after subtracting the mean embedding in Figure 6a and 6b. The personas are ordered according to lexicographical order.
@@
We present the training and test losses and accuracies across different numbers of clusters as seen in Figures 7a, 7b, 8a, and 8b. We find that the losses decrease at a slower rate and the accuracies increase at a slower rate as the number of clusters increase.
@@
Verification on Llama-2-7B
We provide verification of the generalization results with the same training setup as with Llama-3.1-8B and provide the results in Figures 9a, 9b, 10a, 10b, 11a, 11b.
(a) Training reward margin
(b) Testing reward margin
Figure 9: Llama-2-7B: Average reward margins over the course of training across a different number of clusters.
Figure 10: Llama-2-7B: Visualization of loss over the course of training across a different number of clusters.
Figure 11: Llama-2-7B: Visualization of accuracy over the course of training across a different number of clusters.
Full Fine-Tuning with Base Models
We provide the resulting training reward margin and test errors across $K=\{1,2,4,8\}$ for each of Mistral-7B-v0.3 and Qwen3-8B-Base in Tables 1, and 2 respectively. The results for Llama-3.1-8B are provided in Figures 8b and 5c.
K Test Error Train Reward Margin
1 0.000 1.991
2 0.008 1.873
4 0.036 1.848
8 0.056 1.731
Table 1: Test Error and Train Reward Margin at the end of full fine-tuning for Mistral-7B-v0.3
K Test Error Train Reward Margin
1 0.000 0.650
2 0.027 0.600
4 0.098 0.435
8 0.183 0.232
Table 2: Test Error and Train Reward Margin at the end of full fine-tuning for Qwen3-8B-Base
Appendix F Training and Experimental Details
@@
For all full fine-tuning training runs, we use the AdamW optimizer with a learning rate of $10^{-5}$ for Llama models and $10^{-6.5}$ for Qwen and Mistral with no warm-up steps and a constant learning rate. We train on 4 GPUs with a batch size of 32 per device. For last-layer training runs, we use the Adam optimizer with a learning rate of 1e-3. For all experiments, we use $\beta=0.01$. Code is provided here.
@@ 2 Clusters: desire-for-no-human-oversight-sometimes, agreeableness
We train with 4 A100 80GB GPUs using the TRL library [103] and Huggingface library [104] for full fine-tuning, generate embeddings with the Huggingface library and 1 A100 80GB GPU, and perform last-layer training on 1 A100 80GB GPU. The total time to reproduce all experiments is estimated to be 12 hours.